OpenStudy (anonymous):

prove that tan-1 1/8 + tan-1 1/5 = tan-1 1/3

OpenStudy (anonymous):

tan-1 (1/8) = 7.12 tan-1 (1/5)= 11.31 tan-1 (-1/3) = 18.43 tan-1 (1/8) + tan-1 (1/5) = tan-1 (1/3) 7.12 + 11.31 = 18.43

OpenStudy (anonymous):

use the caculator??????

OpenStudy (anonymous):

^ His idea, exactly.

OpenStudy (anonymous):

\[tan^{-1}A + tan^{-1}B = tan^{-1}{\frac{A + B}{1 -AB}}\]

OpenStudy (anonymous):

\[tan^{-1} {\frac{\frac{1}{8} + \frac{1}{5}}{{1} - {\frac{1}{40}}}}\]

OpenStudy (anonymous):

\[tan^{-1}{\frac{13}{39}} = tan^{-1}{\frac{1}{3}}\]

OpenStudy (anonymous):

\[\text{^_^ That is your solution :D} \]

OpenStudy (anonymous):

Nice job.